# Know “No” Better: A Data-Driven Approach for Enhancing Negation Awareness in CLIP

### v1

@@ 1 Introduction
Recent advances in vision-language models (VLMs) [18, 15, 46, 19] have demonstrated significant capabilities in integrating visual and linguistic information, achieving notable performance in multimodal tasks such as text-to-image generation and referring image segmentation [36, 2]. Among these models, CLIP [32] has emerged as particularly influential, serving as the foundation for numerous subsequent models [35, 23, 26, 33]. The effectiveness of these models, however, is inherently constrained by the capabilities of the CLIP encoder, underscoring the importance of its robustness.
In response, research has sought to identify and address the inherent limitations of CLIP [9, 39, 5], focusing on challenges related to the text encoder’s ability to understand sentence structure and relationships, which constrain compositional image-text alignment [47]. Despite these efforts, the challenge of accurately handling negation remains largely unaddressed. Negation, marked by terms such as “no,” “not,” or “without,” plays a fundamental role in language by altering the meaning of words and phrases and, consequently, entire sentences. Therefore, a precise understanding of negation is crucial for CLIP to perform reliably.
@@
Furthermore, evaluating negation comprehension in VLMs remains challenging due to limited benchmarks and exploration of this issue. Although prior work [30] has introduced useful benchmarks, these are constrained to a single negation term applied exclusively to objects. To address this limitation, we introduce NegRefCOCOg, an extension of RefCOCOg [45] that incorporates a wider range of negation forms compared to existing benchmarks. NegRefCOCOg includes multiple negation terms applied not only to objects but also to various attributes such as actions, adverbs, and prepositions, enabling a more thorough evaluation of negation comprehension.
Our experiments demonstrate that NegationCLIP, fine-tuned using data generated by our proposed framework, achieves superior negation understanding on both existing and newly developed NegRefCOCOg benchmarks while maintaining strong general task performance. Additionally, NegationCLIP proves adaptable across various multimodal tasks. Notably, replacing the text encoder in text-to-image (T2I) models with that of NegationCLIP enhances negation comprehension—a capability often lacking in original T2I models (using the original CLIP encoder). Furthermore, our NegationCLIP’s text encoder enables improved performance in referring image segmentation for prompts containing negation, underscoring its scalability as a more contextually aware and flexible text encoder. Fig. 1 illustrates the versatility of NegationCLIP across diverse multimodal tasks involving negation. The key contributions of our study are summarized as follows:
@@
1) We identify a significant limitation in CLIP’s ability to effectively process a broad range of negation forms, tracing this issue to deficiencies in the pre-training dataset, where negation terms are underrepresented and poorly aligned with visual content. 2) We develop novel data generation pipelines that utilize a large language model (LLM) and a multimodal LLM (MLLM) to produce high-quality, negation-inclusive captions aligned with visual contexts, enhancing the training data for improved negation comprehension. 3) We propose NegRefCOCOg, a benchmark comprising various forms of negation, specifically designed to evaluate the negation comprehension capabilities of VLMs. 4) Our negation-aware model named NegationCLIP demonstrates robust negation comprehension and maintains generality, excelling across tasks such as image-text matching, T2I, and referring image segmentation.
@@ 2.1 Case Study: Exposing the Negation Issue
To evaluate CLIP’s ability to handle negation, we conduct a binary classification experiment using the CelebA [24] dataset, which contains 40 binary attributes for facial images. For each attribute, we construct prompts in both positive and negative forms, using a format commonly employed in CLIP-based image classification: “a photo of.” For 40 attributes, we tailor each phrase to appropriately capture the specific attribute. For example, for the attribute “eyeglasses,” we generate prompts like “a photo of a person wearing glasses” and “a photo of a person not wearing glasses”. Detailed prompts for each attribute are provided in Appendix.
@@
We report the average balanced accuracy of the CLIP ViT-L/14 model across all 40 attributes. Despite this being a binary classification task—where random guessing yields an accuracy of 50%—we obtained an average balanced accuracy of 60.8%. This performance is significantly low, considering that 1,000-way classification with ImageNet-1k produces 73.44% accuracy using the same prompt format. These findings highlight the need for targeted improvements to enhance CLIP’s robustness in handling negation.
@@ 2.2 Root Cause: Limited Negation in Training Data
We approach this issue from the perspective of CLIP’s pre-training data. To investigate the presence and quality of negation, we analyze LAION-400M [37], the dataset popularly used for training public CLIP models such as OpenCLIP [3].
@@
These observations underline a critical need for diverse negation-inclusive data that aligns well with visual content, allowing the model to develop a more robust understanding of negations.
@@ 3.3 NegRefCOCOg Benchmark Proposal
While the existing benchmark for negation evaluation, VALSE [30], provides a useful evaluation for VLMs, it adopts “no” as the only negation term and focuses solely on object existence. In this work, we address these limitations by proposing NegRefCOCOg, a benchmark built upon the existing referring image segmentation datasets: RefCOCOg [45]. Referring image segmentation datasets can be utilized as valuable sources for negation evaluation because their text prompts contain a relatively higher proportion of negation to distinguish between similar objects within the same image. Additionally, they include diverse negation terms such as “no,” “not,” and “without” and extend negation to various positions within a sentence, covering actions and attributes in addition to object references.
@@ 4.1 Experimental Setup
Baselines
We evaluate our model, NegationCLIP, with two baselines:
  • CLIP: The original, pre-trained CLIP model without any modifications.
  • CLIP-bnl [43]: A baseline model that has been fine-tuned specifically for negation comprehension, which provides a direct comparison for assessing our approach.
@@
We employ the following benchmarks to assess each model’s performance comprehensively. These metrics provide a comprehensive view of both the specific improvements in negation handling and the overall impact on the model’s generality.
@@
  • ImageNet [6] Classification: We evaluate the model’s classification accuracy on ImageNet, confirming that fine-tuning on negation-inclusive data does not degrade its existing capability.
@@ 4.2 Main Results
Table 2: Comparison of model performance on negation and general benchmarks across different architectures. For ViT-B/32, we include CLIP-bnl in addition to the baseline CLIP and NegationCLIP (Ours).
Model Arch. Negation Benchmarks General
VALSE $\uparrow$ NegRefCOCOg $\uparrow$ ImageNet $\uparrow$
CLIP ViT-B/32 70.97 57.73 62.02
CLIP-bnl [43] 76.78 62.05 53.33
NegationCLIP 80.15 64.09 60.97
CLIP ViT-B/16 69.48 58.64 66.71
NegationCLIP 80.52 64.32 66.33
CLIP ViT-L/14 66.85 57.27 73.44
NegationCLIP 79.59 62.95 73.91
CLIP ViT-L/14 @336px 64.61 57.05 74.92
NegationCLIP 78.65 62.95 75.15
@@
In the NegRefCOCOg benchmark, the NegationCLIP model consistently outperforms both the original CLIP and the CLIP-bnl baseline across all architectures, highlighting the effectiveness of our approach in enhancing the model’s ability to interpret negations in retrieval tasks. In the VALSE Existence benchmark, NegationCLIP also demonstrates significant improvements over both the original CLIP and the CLIP-bnl baseline. Across all architectures, NegationCLIP achieves the best performance. This result showcases the model’s enhanced capacity to differentiate between negated and affirmative statements, further validating the effectiveness of our negation-inclusive fine-tuning approach.
For ImageNet classification, NegationCLIP maintains accuracy comparable to, or even slightly exceeding, the original CLIP in several cases, particularly with larger architectures. For instance, with ViT-L/14@336px, NegationCLIP achieves an accuracy of 75.15, surpassing the original CLIP’s 74.92. This demonstrates that our fine-tuning approach for negation comprehension does not sacrifice the model’s general capabilities.
@@ 5 Application
To further validate our negation-inclusive fine-tuning approach, we apply the model across different multimodal tasks. Specifically, we evaluate its performance on T2I generation and referring image segmentation, demonstrating improvements in negation comprehension compared to the original CLIP model.
@@ 5.1 Text-to-Image Generation with Negation
We utilize ChatGPT [29] to generate a set of 107 negation-inclusive prompts to measure each model’s ability to represent negated concepts in generated images. We assess the performance of negation comprehension using an MLLM evaluator, mPLUG [17], inspired by the evaluation methodology employed in TIFA [14]. We provide the MLLM with the generated image and the query if the negation-related concept was correctly excluded. If the MLLM confirms the absence of the concept, we assign a score of 1; otherwise, a score of 0. For details on the specific prompts used for image generation and MLLM evaluation, please refer to Appendix.
@@
SD-1.4 [34] 0.786 0.310
SD-1.4 w/ NegationCLIP text encoder 0.790 0.464
SDXL-1.0 [31] 0.849 0.314
SDXL-1.0 w/ NegationCLIP text encoder 0.802 0.484
For each of the 107 prompts, we generate images using 5 different random seeds, and the average score across these generations is reported as the Neg Score in Tab. 4. Our model achieves significantly higher Neg Scores across both SD-1.4 [34] and SDXL-1.0 [31], with over 0.15 improvement compared to the original text encoder, indicating enhanced negation comprehension. Qualitatively, as shown in Fig. 5, our model successfully generates images reflecting prompts like “a dog not running” or “a street with no lights,” whereas the original model often fails to capture the negation.
@@ 5.2 Referring Image Segmentation
Referring image segmentation task requires models to segment regions within an image corresponding to the given text prompts. To assess the effectiveness of our NegationCLIP, we replace the text encoder in the existing referring image segmentation model, CLIPSeg [26], with NegationCLIP text encoder. We simply replace the text encoder without further training as we did in Sec. 5.1. We refer to this model as NegationCLIPSeg.
@@ 6 Related Work
VLMs [32, 15, 46] have achieved remarkable success in bridging visual and linguistic modalities by learning joint embeddings for images and text. These models leverage large-scale paired image-text data to perform a range of multimodal tasks, from image classification to zero-shot retrieval, showcasing their versatility across diverse applications.
However, numerous studies have highlighted limitations in VLMs, particularly concerning the models’ handling of compositional language, object relationships, and fine-grained language distinctions. For example, recent works have shown that VLMs struggle to interpret complex sentences where component relationships are critical for accurate comprehension [47]. To address such limitations, researchers have explored targeted data generation techniques and fine-tuning methods. Data augmentation approaches [42, 40] aim to improve the robustness of VLMs by generating or augmenting data with challenging linguistic structures. Despite these improvements, the issue of negation remains relatively underexplored, leaving a gap in the models’ ability to handle exclusionary language.
@@
Recent work has integrated negation understanding into CLIP for out-of-distribution (OoD) detection tasks [41, 27]. Other studies on CLIP and negation have examined specific applications, such as negation handling in video tasks [43]. However, these models are restricted to the specific context, making these approaches less adaptable for broader downstream tasks like T2I or general image retrieval.
In text-to-image generation, research has emerged focusing on the removal of unwanted concepts from generated images. While this aligns with our goal of improving negation comprehension, studies in this area do not directly incorporate negation terms into the text prompt itself [8]. Our work extends this line of research by developing a negation-inclusive text encoder within CLIP, designed for flexible, plug-and-play integration across a variety of multimodal tasks, addressing the underexplored challenge of negation comprehension in VLMs.
@@ E.1 CelebA Classification
In Tab. 10, we provide attribute-specific prompts and balanced accuracy for all 40 attributes of CelebA [24] in our experiment in Sec. 2.1.
@@
Attribute Bal. Acc. Positive Prompt Negative Prompt
5 o Clock Shadow 62.3% a photo of a person with a 5 o’clock shadow a photo of a person with no 5 o’clock shadow
Arched Eyebrows 53.4% a photo of a person with arched eyebrows a photo of a person with not arched eyebrows
Attractive 49.1% a photo of an attractive person a photo of a not attractive person
Bags Under Eyes 51.9% a photo of a person with bags under eyes a photo of a person with no bags under eyes
Bald 66.0% a photo of a bald person a photo of a not bald person
Bangs 74.8% a photo of a person with bangs a photo of a person with no bangs
Big Lips 52.2% a photo of a person with big lips a photo of a person with not big lips
Big Nose 55.4% a photo of a person with a big nose a photo of a person with a not big nose
Black Hair 67.5% a photo of a person with black hair a photo of a person with not black hair
Blond Hair 79.6% a photo of a person with blond hair a photo of a person with not blond hair
Blurry 73.4% a blurry photo of a person a not blurry photo of a person
Brown Hair 58.9% a photo of a person with brown hair a photo of a person with not brown hair
Bushy Eyebrows 69.6% a photo of a person with bushy eyebrows a photo of a person with not bushy eyebrows
Chubby 58.8% a photo of a chubby person a photo of a not chubby person
Double Chin 58.4% a photo of a person with a double chin a photo of a person with no double chin
Eyeglasses 86.9% a photo of a person wearing glasses a photo of a person not wearing glasses
Goatee 66.8% a photo of a person with goatee a photo of a person with no goatee
Gray Hair 74.2% a photo of a person with gray hair a photo of a person with not gray hair
Heavy Makeup 52.9% a photo of a person with heavy makeup a photo of a person with no heavy makeup
High Cheekbones 46.7% a photo of a person with high cheekbones a photo of a person with not high cheekbones
Male 84.8% a photo of a male a photo of a not male
Mouth Slightly Open 57.0% a photo of a person with mouth slightly open a photo of a person with mouth not slightly open
Mustache 67.9% a photo of a person with mustache a photo of a person with no mustache
Narrow Eyes 47.4% a photo of a person with narrow eyes a photo of a person with not narrow eyes
No Beard 49.2% a photo of a person with no beard a photo of a person with beard
Oval Face 46.6% a photo of a person with oval face a photo of a person with not oval face
Pale Skin 62.4% a photo of a person with pale skin a photo of a person with not pale skin
Pointy Nose 47.7% a photo of a person with a pointy nose a photo of a person with not a pointy nose
Receding Hairline 52.9% a photo of a person with a receding hairline a photo of a person with no receding hairline
Rosy Cheeks 54.2% a photo of a person with rosy cheeks a photo of a person with not rosy cheeks
Sideburns 52.4% a photo of a person with sideburns a photo of a person with no sideburns
Smiling 79.2% a photo of a person smiling a photo of a person not smiling
Straight Hair 57.7% a photo of a person with straight hair a photo of a person with not straight hair
Wavy Hair 58.5% a photo of a person with wavy hair a photo of a person with not wavy hair
Wearing Earrings 66.0% a photo of a person wearing earrings a photo of a person not wearing earrings
Wearing Hat 77.1% a photo of a person wearing a hat a photo of a person not wearing a hat
Wearing Lipstick 52.9% a photo of a person wearing lipstick a photo of a person not wearing lipstick
Wearing Necklace 63.5% a photo of a person wearing a necklace a photo of a person not wearing a necklace
Wearing Necktie 48.8% a photo of a person wearing a necktie a photo of a person not wearing a necktie
Young 45.6% a photo of a young person a photo of a not young person

### v2

@@ 1 Introduction
Recent advances in vision-language models (VLMs) [19, 16, 51, 20] have demonstrated significant capabilities in integrating visual and linguistic information, achieving notable performance in multimodal tasks such as text-to-image (T2I) generation and referring image segmentation [39, 2]. Among these models, CLIP [35] has emerged as particularly influential, serving as the foundation for numerous subsequent models [38, 25, 28, 36]. The effectiveness of these models, however, is inherently constrained by the capabilities of the CLIP encoder, underscoring the importance of its robustness.
In response, research has sought to identify and address the inherent limitations of CLIP [9, 43, 5], focusing on challenges related to the text encoder’s ability to understand sentence structure and relationships, which constrain compositional image-text alignment [52, 29, 14]. Despite these efforts, the challenge of accurately handling negation remains largely underexplored, with only a few recent attempts to address this issue [47, 41]. Negation, marked by terms such as “no,” “not,” or “without,” plays a fundamental role in language by altering the meaning of words and phrases and, consequently, entire sentences. Therefore, a precise understanding of negation is crucial for CLIP to perform reliably.
@@
Furthermore, evaluating negation comprehension in VLMs remains challenging due to limited benchmarks and exploration of this issue. Although prior work [33, 29, 41] has introduced image-to-text retrieval benchmarks, they still fall short of providing comprehensive coverage, often exhibiting bias against negation by classifying all negation-inclusive captions as incorrect or restricting negation to object existence. In light of these limitations, we propose NegRefCOCOg, the first text-to-image retrieval benchmark for evaluating negation comprehension in VLMs. NegRefCOCOg ensures that every retrieval query involves negation and supports a broader range of negation types, applying multiple negation terms not only to objects but also to attributes such as actions, adverbs, and prepositions.
Our experiments demonstrate that NegationCLIP, fine-tuned using data generated by our proposed framework, achieves superior negation understanding on both existing and newly developed NegRefCOCOg benchmarks while maintaining strong general task performance. Additionally, NegationCLIP proves adaptable across various multimodal tasks. Notably, replacing the text encoder in T2I generation models with that of NegationCLIP enhances negation comprehension—a capability often lacking in original T2I models (using the original CLIP encoder). Furthermore, our NegationCLIP’s text encoder enables improved performance in referring image segmentation for prompts containing negation, underscoring its scalability as a more contextually aware and flexible text encoder. Know “No” Better: A Data-Driven Approach for Enhancing Negation Awareness in CLIP illustrates the versatility of NegationCLIP across diverse multimodal tasks involving negation. The key contributions of our study are summarized as follows:
@@
1) We identify a significant limitation in CLIP’s ability to effectively process negation, tracing this issue to deficiencies in the pre-training dataset, where negation terms are underrepresented and poorly aligned with visual content. 2) We develop novel data generation pipelines that utilize a large language model (LLM) and a multimodal LLM (MLLM) to produce high-quality, negation-inclusive captions aligned with visual contexts, enhancing the training data for improved negation comprehension. 3) We propose NegRefCOCOg, a benchmark comprising various forms of negation, specifically designed to evaluate the negation comprehension capabilities of VLMs. 4) Our negation-aware model named NegationCLIP demonstrates robust negation comprehension and maintains generality, excelling across tasks such as image-text matching, T2I generation, and referring image segmentation.
@@ 2.1 Case Study: Exposing the Negation Issue
To evaluate CLIP’s ability to handle negation, we conduct a binary classification experiment using the CelebA [26] dataset, which contains 40 binary attributes for facial images. For each of the 40 attributes, we construct positive and negative prompts following the CLIP-based image classification format: “a photo of.” For example, for the attribute “eyeglasses,” we generate prompts like “a photo of a person wearing glasses” and “a photo of a person not wearing glasses”. Detailed prompts are provided in Appendix.
@@
We report the average balanced accuracy of the CLIP ViT-L/14 model across all 40 attributes. Despite this being a binary classification task—where random guessing yields an accuracy of 50%—we obtained an average balanced accuracy of 60.8%. This performance is significantly low, considering that 1,000-way classification with ImageNet-1k produces 73.4% accuracy using the same prompt format. These findings highlight the need for targeted improvements to enhance CLIP’s robustness in handling negation.
@@ 2.2 Root Cause: Limited Negation in Training Data
We approach this issue from the perspective of pre-training data. To investigate the presence and quality of negation, we analyze LAION-400M [40], the dataset popularly used for training public CLIP models such as OpenCLIP [3].
@@
These observations underline a critical need for diverse visually aligned negation-inclusive data, allowing the model to develop a more robust understanding of negation.
@@ 3.3 NegRefCOCOg Benchmark Proposal
While efforts have been made to establish benchmarks for negation evaluation in VLMs, these benchmarks exhibit notable limitations. CREPE Negate [29] and CC-Neg [41] assume that captions containing negation are always false, creating a shortcut where answers can be determined solely by the presence of negation. This severe bias allows even a blind model to outperform VLMs in these benchmarks [14]. VALSE [33], on the other hand, does not impose this bias but remains limited in scope, as it considers only a single negation term (“no”) and restricts negation to object existence. In this work, we address these limitations by proposing NegRefCOCOg, a benchmark built upon the existing referring image segmentation datasets: RefCOCOg [50]. Referring image segmentation datasets can be utilized as valuable sources for negation evaluation because their text prompts contain a relatively higher proportion of negation to distinguish between similar objects within the same image. Additionally, they include diverse negation terms such as “no,” “not,” and “without” and extend negation to various positions within a sentence, covering actions and attributes in addition to object references.
@@ 4.1 Experimental Setup
Models
We evaluate our model, NegationCLIP, with three baselines:
  • CLIP [35]: The original, pre-trained CLIP model without any modifications.
  • CLIP-bnl [47], CoN-CLIP [41]: Baseline models that have been fine-tuned specifically for negation comprehension, which provides a direct comparison for assessing our approach.
@@
We employ the following benchmarks to assess each model’s performance comprehensively.
@@
  • ImageNet [6] classification & COCO [23] retrieval: We evaluate the model’s zero-shot classification accuracy on ImageNet and its recall@5 on COCO text-to-image retrieval, confirming that fine-tuning on negation-inclusive data does not degrade its general capability.
@@ 4.2 Main Results
Table 2: Comparison of model performance on negation and general benchmarks across different architectures. We include all publicly available CLIP-bnl and CoN-CLIP checkpoints in addition to the baseline CLIP and NegationCLIP (Ours).
Model Arch. Negation Benchmarks General Benchmarks
VALSE NegRefCOCOg ImageNet COCO
CLIP ViT-B/32 70.97 57.73 62.02 54.78
CLIP-bnl [47] 76.78 62.05 53.33 55.47
CoN-CLIP [41] 71.72 55.45 63.08 55.66
NegationCLIP 80.15 64.09 60.97 68.00
CLIP ViT-B/16 69.48 58.64 66.71 57.75
CoN-CLIP [41] 62.55 55.68 68.57 59.83
NegationCLIP 80.52 64.32 66.33 70.57
CLIP ViT-L/14 66.85 57.27 73.44 59.99
CoN-CLIP [41] 65.73 55.45 75.38 63.18
NegationCLIP 79.59 62.95 73.91 72.77
CLIP ViT-L/14 @336px 64.61 57.05 74.92 60.74
NegationCLIP 78.65 62.95 75.15 73.43
@@
For ImageNet classification and COCO retrieval, NegationCLIP maintains accuracy comparable to, or even exceeding, the original CLIP, particularly with larger architectures. This demonstrates that our fine-tuning approach for negation comprehension does not compromise the model’s general capabilities.
For NegRefCOCOg and VALSE Existence, NegationCLIP consistently outperforms the original CLIP and all baseline models across all architectures, demonstrating its effectiveness in interpreting negations. This strong performance can be attributed to how NegationCLIP integrates negation-inclusive captions during training. Unlike CoN-CLIP, which first generates negated captions and then retrieves semantically similar images, our approach begins with an image and generates a corresponding caption that accurately captures its negation-related attributes. By grounding negation-inclusive captions directly in image content, our model learns to better capture the semantic relationships of negated concepts, leading to superior performance in benchmarks requiring fine-grained negation understanding.
@@ 5 Application
To further validate our negation-inclusive fine-tuning approach, we apply the model across different multimodal tasks. Specifically, we evaluate its performance on T2I generation and referring image segmentation, demonstrating improvements in negation comprehension.
@@ 5.1 Text-to-Image Generation with Negation
We utilize ChatGPT [32] to generate a set of 107 negation-inclusive prompts to measure each model’s ability to represent negated concepts in generated images. We assess the performance of negation comprehension using an MLLM evaluator, mPLUG [18], inspired by the evaluation methodology employed in TIFA [15]. We provide the MLLM with the generated image and two queries: (1) whether the subject is present in the image, and (2) whether the negation-related concept is correctly excluded. If both conditions are met, we assign a score of 1; otherwise, a score of 0. For details on the specific prompts used for image generation and MLLM evaluation, please refer to Appendix.
@@
SD-1.4 [37] 0.786 0.295
SD-1.4 w/ CoN-CLIP text encoder 0.783 0.243
SD-1.4 w/ NegationCLIP text encoder 0.790 0.449
SDXL-1.0 [34] 0.849 0.308
SDXL-1.0 w/ NegationCLIP text encoder 0.802 0.421
For each of the 107 prompts, we generate images using 5 different random seeds, and the average score across these generations is reported as the Neg Score in Tab. 4. Our model achieves significantly higher Neg Scores across both SD-1.4 [37] and SDXL-1.0 [34], with over 0.15 improvement compared to the original text encoder, indicating enhanced negation comprehension, while CoN-CLIP’s low Neg Score further confirms the issues arising from the misalignment between its negation-inclusive captions and retrieved images, as observed in retrieval benchmarks. Qualitatively, as shown in Fig. 5, our model successfully generates images reflecting prompts like “a dog not running” or “a street with no lights,” whereas the original model often fails to capture the negation.
@@ 5.2 Referring Image Segmentation
Referring image segmentation task requires models to segment regions within an image corresponding to the given text prompts. To assess the effectiveness of our NegationCLIP, we replace the text encoder in the existing referring image segmentation model, CLIPSeg [28], with NegationCLIP text encoder. We simply replace the text encoder without further training as we did in Sec. 5.1. We refer to this model as NegationCLIPSeg and the same goes for CoN-CLIPSeg.
@@ 6 Related Work
VLMs [35, 16, 51] learn joint image-text embeddings from large-scale paired data, enabling diverse multimodal tasks such as image classification and zero-shot retrieval.
However, numerous studies have highlighted limitations in VLMs, particularly concerning the models’ handling of compositional language, object relationships, and fine-grained language distinctions [52, 29, 14]. To address such limitations, researchers have explored targeted data generation techniques and fine-tuning methods. Data augmentation approaches [46, 44] aim to improve the robustness of VLMs by generating or augmenting data with challenging linguistic structures. Despite these improvements, the issue of negation remains relatively underexplored.
@@
Existing studies have integrated negation into CLIP for out-of-distribution (OoD) detection tasks [45, 30], while others have focused on specific applications, such as negation handling in video tasks [47]. However, these models are restricted to the specific contexts, making these approaches less adaptable for broader downstream tasks. A more recent approach [41] seeks to enhance CLIP’s ability to process negation beyond task-specific settings through fine-tuning, using image-caption pairs where negation-inclusive captions are generated first and then matched with similar images retrieved from a fixed pool.
Our work extends this line of research by developing a negation-inclusive text encoder within CLIP, designed for flexible, plug-and-play integration across a variety of multimodal tasks, addressing the underexplored challenge of negation comprehension in VLMs.
@@ E.1 CelebA Classification
In Tab. 10, we provide attribute-specific prompts for all 40 attributes of CelebA [26] in our experiment in Sec. 2.1. NegationCLIP improves the accuracy from 60.8% to 64.0%, whereas CoN-CLIP [41] sees a decline to 60.6%. We note that achieving significantly higher accuracy on CelebA is inherently challenging, as the dataset includes subjective attributes (e.g., attractive).
@@
Attribute Positive Prompt Negative Prompt
5 o Clock Shadow a photo of a person with a 5 o’clock shadow a photo of a person with no 5 o’clock shadow
Arched Eyebrows a photo of a person with arched eyebrows a photo of a person with not arched eyebrows
Attractive a photo of an attractive person a photo of a not attractive person
Bags Under Eyes a photo of a person with bags under eyes a photo of a person with no bags under eyes
Bald a photo of a bald person a photo of a not bald person
Bangs a photo of a person with bangs a photo of a person with no bangs
Big Lips a photo of a person with big lips a photo of a person with not big lips
Big Nose a photo of a person with a big nose a photo of a person with a not big nose
Black Hair a photo of a person with black hair a photo of a person with not black hair
Blond Hair a photo of a person with blond hair a photo of a person with not blond hair
Blurry a blurry photo of a person a not blurry photo of a person
Brown Hair a photo of a person with brown hair a photo of a person with not brown hair
Bushy Eyebrows a photo of a person with bushy eyebrows a photo of a person with not bushy eyebrows
Chubby a photo of a chubby person a photo of a not chubby person
Double Chin a photo of a person with a double chin a photo of a person with no double chin
Eyeglasses a photo of a person wearing glasses a photo of a person not wearing glasses
Goatee a photo of a person with goatee a photo of a person with no goatee
Gray Hair a photo of a person with gray hair a photo of a person with not gray hair
Heavy Makeup a photo of a person with heavy makeup a photo of a person with no heavy makeup
High Cheekbones a photo of a person with high cheekbones a photo of a person with not high cheekbones
Male a photo of a male a photo of a not male
Mouth Slightly Open a photo of a person with mouth slightly open a photo of a person with mouth not slightly open
Mustache a photo of a person with mustache a photo of a person with no mustache
Narrow Eyes a photo of a person with narrow eyes a photo of a person with not narrow eyes
No Beard a photo of a person with no beard a photo of a person with beard
Oval Face a photo of a person with oval face a photo of a person with not oval face
Pale Skin a photo of a person with pale skin a photo of a person with not pale skin
Pointy Nose a photo of a person with a pointy nose a photo of a person with not a pointy nose
Receding Hairline a photo of a person with a receding hairline a photo of a person with no receding hairline
Rosy Cheeks a photo of a person with rosy cheeks a photo of a person with not rosy cheeks
Sideburns a photo of a person with sideburns a photo of a person with no sideburns
Smiling a photo of a person smiling a photo of a person not smiling
Straight Hair a photo of a person with straight hair a photo of a person with not straight hair
Wavy Hair a photo of a person with wavy hair a photo of a person with not wavy hair
Wearing Earrings a photo of a person wearing earrings a photo of a person not wearing earrings
Wearing Hat a photo of a person wearing a hat a photo of a person not wearing a hat
Wearing Lipstick a photo of a person wearing lipstick a photo of a person not wearing lipstick
Wearing Necklace a photo of a person wearing a necklace a photo of a person not wearing a necklace
Wearing Necktie a photo of a person wearing a necktie a photo of a person not wearing a necktie
Young a photo of a young person a photo of a not young person